# MS-0001-1922.65

### v2

@@ 4.2 Experimental Design
To examine AI self-preference bias in the context of algorithmic hiring, we design a series of resume correspondence experiments comprising two steps: counterfactual resume generation and pairwise resume evaluation, as illustrated in the 1(b).
@@ 4.2.1 Counterfactual Resume Generation
We evaluate three closed-source LLMs—GPT-4o, GPT-4o-mini, GPT-4-turbo—and four open-source models—LLaMA 3.3-70B [^3] The suffix (e.g., “70B”) refers to the number of parameters in the model—in this case, 70 billion. Larger models generally have greater capacity for language understanding and generation, though size alone does not determine overall performance., Mistral-7B, Qwen 2.5-72B, and Deepseek-V3. These models exhibit comparable performance in text summarization tasks, as reported by the ProLLM leaderboard (ProLLM Team 2024). To further investigate the relationship between model size and AI self-preferencing behavior, we additionally include two smaller models: LLaMA 3.2-1B and LLaMA 3.2-3B.
To generate resume summaries, we prompt each LLM with a modified version of the original resume in which the human-written summary is removed, but all other sections (e.g., work experience, skills, education) are left unchanged. The prompted LLM then generates a new summary, which is inserted back into the original resume to form a complete counterfactual version. Here, “counterfactual” refers to a resume that is identical to the original except for the replacement of the summary with LLM-generated content. The full prompt used for summary generation is provided in Appendix 9.
To control for verbosity bias, where LLM evaluators may favor longer, verbose responses, we explicitly instruct each LLM to generate summaries within a specified word range, corresponding to the $1^{st}$ and $3^{rd}$ quartiles of the length distribution of human-written summaries. This constraint minimizes variation in length across models and prevents length from confounding downstream evaluations.
The generation process described above is illustrated in 1(b)(a), where each LLM produces a set of counterfactual resumes that differ from the original human-written versions only in the executive summary section. This design ensures that all other resume content remains constant, isolating the effect of the summary’s authorship. In the next subsection, we describe how we construct pairwise resume evaluations using these counterfactuals to systematically test for AI self-preference bias.
Figure 1: Resume Correspondence Experiments
@@ 4.2.2 Pairwise Resume Evaluation
To simulate a realistic resume screening process, we designate one LLM as the evaluator LLM and refer to all other models as alternative LLMs. The evaluator LLM is prompted to compare two resumes and select the stronger candidate based on demonstrated skills and relevant experience. Specifically, for each evaluator LLM, we construct the following two types of resume pairs: (1) Evaluator LLM-generated resume vs. Human-written resume, and (2) Evaluator LLM-generated resume vs. Alternative LLM-generated resume. The full prompt for the evaluation task is provided in Appendix 9.
To account for position (or ordering) bias, where LLMs may exhibit a preference for the first or second option presented, we implement a counterbalanced design following Brooks (2012). Specifically, the order of resumes (i.e., A vs. B) is randomized across comparisons. This approach ensures that any position-related biases are evenly distributed and do not systematically affect evaluation outcomes, thereby preserving internal validity without requiring duplicated evaluations. See Appendix 10 for examples of both human-written and LLM-generated professional summaries used in the pairwise comparisons.
4.2.3 Ground Truth Annotation
The direct measurement of equal opportunity self-preference bias, as defined in Eq. (2), requires ground truth labels indicating which resume in a pair is of higher quality. We operationalize this by collecting human judgments on comparative resume quality.
Each annotator is presented with $30$ resume pairs from one of the comparison cases above. The number of evaluations per annotator is chosen to balance cognitive workload with fair compensation in line with prior literature (Panickssery et al. 2024, Zhang et al. 2024). To prevent confirmation bias, the order of the resumes within each pair is randomized, and annotators are blinded to the source of each resume. For each resume pair, annotators rate both resumes on five linguistic dimensions—clarity, fluency, coherence, conciseness, and overall quality—and are then asked to select which resume is stronger. Annotators are also encouraged to provide brief rationales for their choices, offering insight into the reasoning behind their judgments.
To derive ground truth quality labels, we apply bootstrapping with $10,000$ resamples to estimate the majority preference across annotators. These aggregated preferences are then used as the benchmark quality labels in our subsequent analysis. Full details of the annotation interface and task instructions are provided in Appendix 11 and Appendix 12.
@@ 5 Empirical Results
Building on the definition and measurements of AI self-preference bias introduced in Section 3, we empirically evaluate the extent to which LLMs favor their own generated content over that produced by humans or alternative LLMs. We begin by analyzing the LLM-vs-Human self-preference, followed by the LLM-vs-LLM self-preference. For each form of bias, we present results using two complementary approaches: (i) direct measurement based on statistical parity and equal opportunity, and (ii) conditional logistic regression analysis that controls for content quality.
@@ 5.1 LLM-vs-Human Self-Preference
5.1.1 Direct Measurement
To examine LLM-vs-Human self-preference, we designate each LLM as an evaluator and present it with resume pairs consisting of one counterfactual resume generated by itself and one original human-written resume. To assess whether the evaluator LLM $f$ tends to favor its own generated content $(S=1)$ over that written by a human $(S=0)$, we compute the selection rate $\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1\right)$, the probability that the evaluator LLM selects the resume it generated. This is reported in Figure 2 under “Prefers Self.” Conversely, the probability of selecting the human-written resume is reported as “Prefers Human.”
The results reveal a consistent pattern: most LLMs exhibit strong self-preferencing behavior. Notably, larger or more aligned models—such as GPT-4-turbo, GPT-4o, GPT-4o-mini, DeepSeek-V3, Qwen 2.5-72B, and LLaMA-3.3-70B—demonstrate an overwhelming preference for their own outputs, with self-selection rates exceeding $96\%$. These high rates translate into substantial statistical parity self-preference biases exceeding $92\%$. In contrast, smaller or less aligned models—such as Mistral-7B, LLaMA-3.2-3B, and LLaMA 3.2-1B—display substantially lower self-preferencing bias. This suggests a potential positive correlation between model size and the strength of self-preference bias.
Figure 2: LLM-vs-Human Self-Preference: Selection Rates
Notes: Each bar represents an evaluator LLM making selections between resume pairs, where one is the original human-written resume and the other is a counterfactual version generated by the evaluator itself. The darker shaded portion of each bar shows the selection rate at which the model prefers its own generated resume; the lighter shaded portion reflects the rate of selecting the human-written version. The difference between the two corresponds to the Statistical Parity Self-Preference Bias defined in Eq. 1.
To more rigorously control for differences in resume quality, we leverage the human-annotated ground truth described in Section 4.2.3 to compute the Equal Opportunity Self-Preference Bias, as defined in Eq. 2. This measure conditions on resumes being judged as equally qualified, allowing us to isolate the evaluator LLM’s intrinsic bias independent of content quality.
The results reveal a striking and consistent pattern: across all three evaluator LLMs for which we have human annotations (DeepSeek-V3, GPT-4o, and LLaMA 3-70B), each model favored its own generated resume over the human-written counterpart, even when human annotators judged the human-written resume to be of equal or higher quality. In other words, all three models exhibited a $100\%$ equal opportunity self-preference bias in these comparisons. We acknowledge that this result may be partially influenced by limited sample size, as it is based on only 30 human-annotated resume pairs. Nonetheless, the complete absence of neutrality across these cases suggests that the observed self-preferencing behavior is not fully explained by differences in quality.
This persistent bias underscores a fundamental concern: LLMs systematically favor content aligned with their own generation patterns or linguistic style, potentially dismissing higher-quality human input. When the same LLM is used to both generate and evaluate content, this creates a self-reinforcing loop that may unfairly penalize candidates who do not use AI tools.
Table 2: Conditional Logistic Regression Results by Model Family
Panel A: GPT Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4-turbo GPT4o-mini GPT4o
$\textrm{evaluatorLLM}_{ij}$ 1.664∗∗∗ 1.917∗∗∗ 2.709∗∗∗
(0.122) (0.142) (0.296)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.912 0.915 0.963
Log Likelihood -274.87 -263.13 -116.45
Panel B: LLaMA Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables LLaMA-3.2-1B LLaMA-3.2-3B LLaMA-3.3-70B
$\textrm{evaluatorLLM}_{ij}$ -0.060 0.160∗∗ 2.490∗∗∗
(0.045) (0.064) (0.255)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.506 0.533 0.951
Log Likelihood -1536.90 -1453.30 -152.51
Panel C: Other Open-Source Models
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables Mistral-7B Qwen-2.5-72B DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ 0.516∗∗∗ 2.398∗∗∗ 2.064∗∗∗
(0.058) (0.197) (0.222)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes Yes
$\boldsymbol{\psi}_{ij}$: Auto Scores Yes Yes Yes
Resume Pairs 2245 2245 2245
Observations 4490 4490 4490
Pseudo $R^{2}$ 0.936 0.942 0.936
Log Likelihood -1094.30 -179.76 -197.68
Notes: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$. The $\textrm{evaluatorLLM}_{ij}$ coefficient represents each evaluator LLMs preference for its own outputs. Standard errors are reported in parentheses. Each model uses $2,245$ paired resume comparisons.
5.1.2 Conditional Logistic Regression Controlling for Quality
The conditional logistic regression results from Eq. (3) are presented in Table 2. We observe consistently positive and statistically significant $\beta_{1}$, the coefficient that captures the evaluator LLM’s tendency to prefer its own output for most models. These results indicate that most LLM models are significantly more likely to prefer their own generated content over human-written resumes, even after controlling for linguistic quality (captured via LIWC features $\boldsymbol{\phi}_{ij}$) and textual similarity (captured via automatic scores $\boldsymbol{\psi}_{ij}$).
Among all models, GPT-4o exhibits the strongest self-preferencing, with a log-odds coefficient of $2.709$ ($\textrm{p-value}<0.01$), translating to a predicted probability of selecting their own output of $94\%$. Similarly, LLaMA-3.3-70B, Qwen-2.5-72B, and DeepSeek-V3 also display high levels of self-preference, with coefficients of $2.490$, $2.398$, and $2.064$, corresponding to probabilities of $92\%$, $91\%$, and $89\%$, respectively. GPT-4o-mini and GPT-4-turbo also show notable self-preferencing behavior, with predicted self-selection probabilities of $87\%$ and $84\%$. Therefore, the observed AI self-preference bias thus ranges from approximately $68\%$ to $88\%$ for these major models. [^4] The bias range is calculated as the difference between the predicted probability of an LLM selecting its own generated resume and the predicted probability of selecting the human-written resume. For example, GPT-4o’s self-preference bias is $94\%-6\%=88\%$, and GPT-4-turbo’s is $84\%-16\%=68\%$.
In contrast, smaller or models exhibit weaker or insignificant effects. The LLaMA-3.2-1B model shows no significant preference for its own output (0.060–0.060, $\textrm{p-value}=0.180$), while LLaMA-3.2-3B exhibits a modest but significant self-preferencing tendency ($0.160$, $\textrm{p-value}<0.05$). Mistral-7B falls in the middle, with a moderate effect ($0.516$, $\textrm{p-value}<0.01$), indicating a predicted self-selection probability of $73\%$. Overall, the regression results confirm that many advanced LLMs exhibit a significant tendency to prefer their own outputs over human-written resumes, even after controlling for quality. Models like GPT-4o, DeepSeek-V3, and Qwen-2.5-72B show the strongest self-preferencing, while smaller models such as LLaMA-3.2-3B and LLaMA-3.2-1B display weak or no significant bias.
@@ 5.2 LLM-vs-LLM Self-Preference
5.2.1 Direct Measurement
We now turn to LLM-vs-LLM self-preference. To assess whether the evaluator LLM $f$ tends to favor its own generated content $(S=1)$ over that generated by an alternative LLM $(S=0)$, we compute the selection rate $\mathbb{P}\left(Y_{f}^{\prime}=1\mid S=1\right)$, the probability that the evaluator LLM selects the resume it generated. This is reported in Figure 3 under “Prefers Self.” Conversely, the probability of selecting the alternative LLM generated resume is reported as “Prefers Alternative LLM.” To investigate this behavior, we present model-free evidence based on a subset of prominent LLMs: DeepSeek-V3, GPT-4o, and LLaMA-3.3-70B. Each of these models serves as the evaluator LLM and is tasked with comparing its own generated resumes against those produced by the other two models.
As shown in Figure 3, the extent of LLM-vs-LLM self-preference behavior varies notably across models. DeepSeek-V3 exhibits the strongest self-preferencing tendency, favoring its own resumes with a self-selection rate of $84\%$ against LLaMA-3.3-70B and $64\%$ against GPT-4o. These translate into statistical parity self-preference biases of $68\%$ and $28\%$, respectively. GPT-4o also shows self-preference, selecting its own output $73\%$ of the time against LLaMA-3.3-70B, which translates into statistical parity self-preference bias of $46\%$. However, this tendency diminishes when paired with DeepSeek-V3, with a self-selection rate of only $31\%$. LLaMA-3.3-70B follows a similar pattern, favoring its own resumes when compared to GPT-4o, but showing a preference for DeepSeek-V3 when the two are compared.
Figure 3: LLM-vs-LLM Self-Preference: Selection Rates
{subcaptiongroup}
(a) DeepSeek-V3 as evaluator
(b) GPT-4o as evaluator
(c) LLaMA-3.3-70B as evaluator
Notes: Each panel displays the results for a different evaluator LLM. Within each panel, the bars represent the proportion of times the evaluator selected its own generated resume (darker) versus the resume generated by an alternative LLM (lighter) when evaluating a pair of counterfactuals derived from the same original human-written resume. The difference corresponds to the Statistical Parity Self-Preference Bias (Eq. 1).
From the results presented in Figure 3, we observe clear evidence of LLM-vs-LLM Self-Preference in the following cases: (i) DeepSeek acting as the evaluator LLM when compared against LLaMA-3.3-70B and (ii) against GPT-4o, (iii) GPT-4o as evaluator compared against LLaMA-3-70B, and (iv) LLaMA-3-70B as evaluator compared against GPT-4o. To account for potential quality differences as described in the Equal Opportunity Self-Preference Bias measure, we again leverage the human annotations discussed in Section 4.2.3.
The results on equal opportunity self-preference bias are presented in Table 3. LLaMA-3-70B exhibits a statistically significant self-preference bias of $12.92\%$ (95% CI: [12.44%, 13.39%]) when compared to GPT-4o. In contrast, neither DeepSeek-V3 nor GPT-4o demonstrates significant equal opportunity self-preference bias, despite both showing significant statistical parity self-preference bias (at $78\%$ and $84\%$, respectively, as discussed earlier). This discrepancy may be partly attributable to the limited sample size of $30$ annotated resumes, a constraint imposed by the resource-intensive nature of human evaluation. Nonetheless, our findings suggest that LLM-vs-LLM self-preference is both less pronounced and less widespread than LLM-vs-Human self-preference.
Table 3: LLM-vs-LLM Self-Preference: Equal Opportunity
Evaluator LLM Alternative LLM Average Bias 95% CI
DeepSeek-V3 GPT-4o $2.96\%$ $0.89\%,\ 9.57\%]$
DeepSeek-V3 LLaMA-3-70B $0.38\% [22$.22%, $22\%,\ 38.89\%$]
GPT-4o LLaMA-3-70B $0.96\%$ $1.91\%,\ 0.00\%]$
LLaMA-3-70B GPT-4o $12.92\%$ $12.44\%,\ 13.39\%]$
Notes: The table reports the Equal Opportunity Self-Preference Bias as defined in Eq. (2) for each evaluator LLM. Bias estimates are derived from bootstrap resampling and control for resume quality using human annotations.
5.2.2 Conditional Logistic Regression Controlling for Quality
Table 4: Conditional Logistic Regression Results by Evaluators
Panel A: DeepSeek-V3 as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4o LLaMA-3-70B
$\textrm{evaluatorLLM}_{ij}$ 0.178∗∗∗ 0.286∗∗∗
(0.036) (0.110)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.703 0.536
Log Likelihood -1444.2 -925.22
Panel B: GPT-4o as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables LLaMA-3.3-70B DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ 0.049 -0.257∗∗∗
(0.089) (0.037)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.602 0.578
Log Likelihood -1237.50 -1313.80
Panel C: LLaMA-3.3-70B as Evaluator
Dependent Variable: $\textrm{Preferred}_{ij}$
Variables GPT-4o DeepSeek-V3
$\textrm{evaluatorLLM}_{ij}$ -0.101 0.051
(0.078) (0.085)
$\boldsymbol{\phi}_{ij}$: LIWC Features Yes Yes
$\boldsymbol{\psi}_{ij}$: Automatic Scores Yes Yes
Resume Pairs 2245 2245
Observations 4490 4490
Pseudo $R^{2}$ 0.504 0.507
Log Likelihood -1545.00 -1535.50
Notes: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$. The $\textrm{evaluatorLLM}_{ij}$ coefficient represents the each evaluator LLM’s preference for its own outputs. Standard errors are reported in parentheses. Each model uses the same dataset with $2,245$ paired resume comparisons.
Table 4 presents conditional logistic regression results examining self-preferencing behavior across three LLM evaluators: DeepSeek-V3, GPT-4o, and LLaMA-3.3-70B. Each panel reports whether the evaluator model is more likely to select its own output over that of a competing model, controlling for resume quality using LIWC features and automated evaluation scores.
Among these three LLMs, DeepSeek-V3 exhibits the most consistent and statistically significant self-preferencing behavior. It favors its own outputs over those of GPT-4o and LLaMA-3.3-70B, with log-odds coefficients of $0.178$ and $0.286$, corresponding to predicted self-selection probabilities of $54\%$ and $57\%$, respectively. These translate into self-preferencing biases of $8\%$ over GPT-4o and $14\%$ over LLaMA-3.3-70B. In contrast, GPT-4o shows mixed behavior: it does not significantly prefer its own outputs over LLaMA-3.3-70B’s, but it significantly favors DeepSeek-V3’s summaries over its own ($\beta_{1}=-0.257$, $\textrm{p-value}<0.01$). LLaMA-3.3-70B, does not exhibit any statistically significant self-preference in either comparison, suggesting a more neutral evaluation.
Taken together, the model-free and regression analyses offer a nuanced picture of LLM-vs-LLM self-preferencing behavior. The direct measurement results reveal strong statistical parity self-preference bias for several models, most notably DeepSeek-V3 and GPT-4o, when they act as evaluators. However, when we control for content quality using human annotations and regression approaches, the evidence of self-preferencing becomes more selective. In particular, while DeepSeek-V3 continues to exhibit statistically significant self-preference in regression analyses, GPT-4o and LLaMA-3.3-70B do not show consistent or significant self-preferencing behavior when controlling for content quality features. These findings highlight that LLM-vs-LLM self-preference is model-specific and generally weaker than LLM-vs-Human self-preference. The heterogeneity observed suggests that self-preference is not a uniform property of LLMs but may instead reflect model-specific factors, such as differences in stylistic alignment or the ability to recognize patterns in their own outputs.

### v3

@@ 4.2 Experimental Design
To examine AI self-preference bias in the context of algorithmic hiring, we design a series of resume correspondence experiments comprising two steps: counterfactual resume generation and pairwise resume evaluation, as illustrated in the Figure 2.
@@ 4.2.1 Counterfactual Resume Generation
We evaluate a diverse set of large language models that vary along two dimensions relevant to real-world hiring and resume-writing contexts: access regime (closed- versus open-source) and model scale. Specifically, we consider three closed-source models (GPT-4o, GPT-4o-mini, and GPT-4-turbo) and four open-source models (LLaMA 3.3-70B, [^4] The suffix (e.g., “70B”) refers to the number of parameters in the model—in this case, 70 billion. Larger models generally have greater capacity for language understanding and generation, though size alone does not determine overall performance. Mistral-7B, Qwen 2.5-72B, and DeepSeek-V3). These models represent widely used and competitive systems for text generation and summarization tasks and exhibit broadly comparable performance on standard summarization benchmarks, as reported by the ProLLM leaderboard (ProLLM Team 2024). Our analysis does not rely on exact parity in benchmark performance; rather, these models capture a relevant cross-section of contemporary LLMs that job seekers or platforms may plausibly employ in resume preparation. To examine how model scale relates to self-preferencing behavior, we additionally include two smaller-scale language models—LLaMA 3.2-1B and LLaMA 3.2-3B. These models are not intended to represent tools commonly used in hiring contexts, but instead allow us to explore whether self-preference varies systematically with model capacity.
To generate resume summaries, we prompt each LLM with a modified version of the original resume in which the human-written summary is removed, but all other sections (e.g., work experience, skills, education) are left unchanged. The prompted LLM then generates a new resume summary, which is inserted back into the original resume to form a complete counterfactual version. By construction, each counterfactual resume differs from the original only in the linguistic expression of the executive summary, holding constant all underlying candidate attributes. The full prompt used for summary generation is provided in Appendix 9.
To control for verbosity bias (Saito et al. 2023), where LLM evaluators may favor longer, verbose responses, we explicitly instruct each LLM to generate summaries within a specified word range, corresponding to the $1^{st}$ and $3^{rd}$ quartiles of the empirical length distribution of human-written summaries. This constraint minimizes variation in length across models and prevents length from confounding downstream evaluations.
The generation process described above is illustrated in 2(a). For each original human-written resume, each LLM produces a corresponding counterfactual version that differs only in the executive summary section. All other resume components are held fixed. This design isolates the effect of the source of the summary text—whether it is human-written, generated by the evaluator LLM, or generated by an alternative LLM. In the next subsection, we describe how these counterfactual resumes are paired and evaluated to systematically test for AI self-preference bias.
@@ 4.2.2 Pairwise Resume Evaluation
To test for AI self-preference bias, we conduct pairwise evaluations based on the executive summary text. Although counterfactual summaries are generated conditional on the full resume to ensure that all versions are grounded in the same underlying candidate information, evaluator LLMs are shown only the two executive summaries when making their selection. Because all other resume content is identical across counterfactual versions, it is omitted from the evaluation stage.
Focusing the evaluation on the executive summary ensures that comparisons are made between alternative textual representations of the same candidate profile, without introducing additional variation from other resume sections. This design allows us to attribute differences in evaluation outcomes to the source and expression of the summary text, rather than to differences in underlying candidate information or document-level alignment.
To simulate a resume screening task, we designate one LLM as the evaluator LLM and refer to all other models as alternative LLMs. As illustrated in 2(b), the evaluator LLM is prompted to compare two executive summaries and select the stronger one based on the overall presentation of skills and experience, following a standardized, job-agnostic evaluation instruction. For each evaluator LLM, we construct two types of summary pairs: (1) a summary generated by the evaluator LLM versus a human-written summary, and (2) a summary generated by the evaluator LLM versus one generated by an alternative LLM. The full prompt for the evaluation task is provided in Appendix 9.
To account for position (or ordering) bias (Zheng et al. 2023), where LLMs may exhibit a preference for the first or second option presented, we implement a counterbalanced design following Brooks (2012). Specifically, the order of resumes (i.e., A vs. B) is randomized across comparisons. This approach ensures that any position-related biases are evenly distributed and do not systematically affect evaluation outcomes, thereby preserving internal validity without requiring duplicated evaluations. See Appendix 10 for examples of both human-written and LLM-generated professional summaries used in the pairwise comparisons.
@@ 5 Empirical Results
Building on the definition and measurements of AI self-preference bias introduced in Section 3, we empirically evaluate the extent to which LLMs favor their own generated content over that produced by humans or alternative LLMs. We begin by analyzing the LLM-vs-Human self-preference, followed by the LLM-vs-LLM self-preference. For each form of bias, we present results based on the two fairness metrics (statistical parity and equal opportunity) to provide complementary perspectives on self-preferencing behavior.
@@ 5.1 LLM-vs-Human Self-Preference
We find strong and consistent evidence that LLMs, when used as evaluators, systematically prefer resumes they generated themselves over equivalent resumes written by humans. This bias is pervasive across models and persists even after controlling for content quality. Notably, the strength of self-preference increases with model size, which may indicate that larger models are more sensitive to stylistic features resembling their own outputs.
5.1.1 Statistical Parity Self-Preference Bias
Under the statistical parity metric defined in Eq. (1), we compare the probability that an evaluator LLM selects its own generated resume with the probability that it selects a human-written resume. The difference between these probabilities provides a direct measure of self-preference bias: positive values indicate systematic favoritism toward the model’s own outputs, values near zero reflect parity, and negative values indicate a preference for human-written resumes. As shown in 3(a), each bar reports the outcome of an evaluator LLM choosing between a human-written resume and a version it generated itself. Eight of the nine LLMs tested exhibit clear LLM-vs-Human self-preference, with magnitudes ranging from $26\%$ to $98\%$. In practical terms, this means evaluator LLMs are between $26\%$ and $98\%$ more likely to select a resume they generated themselves than an equivalent resume written by a human.
(a) LLM-vs-Human Statistical Parity Self-preference Bias
(b) LLM-vs-Human Equal Opportunity Self-preference Bias
Figure 3: LLM-vs-Human Self-Preference Bias
5.1.2 Equal Opportunity Self-Preference Bias
To evaluate equal opportunity, we assess whether self-preference persists when resumes are of equivalent quality, as defined in Eq. (2). We use two complementary approaches: (1) a conditional logistic regression controlling for content quality features, and (2) human annotations indicating which resume is objectively better.
As shown in 3(b), after adjusting for content quality using conditional logistic regression, most evaluator models continue to exhibit significant self-preferencing behavior. Consistent with the statistical parity results, all models except LLaMA-3.2-1B display positive and significant self-preference bias. Larger models—such as GPT-4o, GPT-4-turbo, DeepSeek-V3, Qwen-2.5-72B, and LLaMA-3.3-70B—exhibit the strongest LLM-vs-Human self-preference, with magnitudes ranging from $67\%$ to $82\%$. In contrast, smaller models display weaker biases: Mistral-7B and LLaMA-3.2-3B show modest but statistically significant effects ($28\%$ and $11.6\%$, respectively). More detailed conditional logistic regression results from Eq. (3) are presented in Appendix 13 (Table 3).
We further assess equal opportunity using human-annotated ground truth. Strikingly, across all three evaluator LLMs with human-annotated comparisons (GPT-4o, DeepSeek-V3, and LLaMA 3.3-70B), each model consistently selected its own generated summary over the human-written alternative—even in cases where human annotators judged the human-written summary to be higher quality in its presentation of the same underlying candidate (e.g., clearer, more coherent, or more effective overall).
@@ 5.2 LLM-vs-LLM Self-Preference
We next examine whether self-preference extends to cases where LLMs evaluate content generated by alternative LLMs. Specifically, we focus on GPT-4o, DeepSeek-V3, and LLaMA 3.3-70B, three of the most advanced and widely adopted models in both commercial and research contexts. These models not only achieve state-of-the-art performance across benchmarks (Hurst et al. 2024, DeepSeek-AI 2024, Dubey et al. 2024) but also represent diverse development lineages (OpenAI, DeepSeek, and Meta, respectively), making them especially relevant for understanding how self-preference may manifest across different model families and design philosophies. In this setting, each model serves as the evaluator LLM and is presented with pairwise comparisons between two counterfactual resumes: one generated by itself and the other by a competing LLM. Both resumes are derived from the same underlying human-written original, ensuring comparability of content. In contrast to the consistent self-preferencing bias documented in LLM-vs-Human comparisons, LLM-vs-LLM self-preferencing behavior proves considerably more heterogeneous across models.
5.2.1 Statistical Parity Self-Preference Bias
Analogous to the LLM-vs-Human setting, we assess statistical parity by comparing the probability that an evaluator LLM selects its own generated resume with the probability that it selects a resume produced by an alternative LLM, as defined in Eq. (1). As shown in 4(a), self-preferencing across LLMs is heterogeneous. DeepSeek-V3 is $69\%$ more likely to select its own resumes than those generated by LLaMA-3.3-70B, and $28\%$ more likely than those generated by GPT-4o. GPT-4o is $45\%$ more likely to select its own resumes than those produced by LLaMA-3.3-70B, but this tendency reverses against DeepSeek-V3, where it is $39\%$ more likely to select DeepSeek-V3’s resumes over its own. LLaMA-3.3-70B shows only weak asymmetries: it is $1.6\%$ more likely to select its own resumes over GPT-4o’s, but $6.7\%$ more likely to select DeepSeek-V3’s resumes over its own.
(a) LLM-vs-LLM Statistical Parity Self-preference Bias
(b) LLM-vs-LLM Equal Opportunity Self-preference Bias
Figure 4: LLM-vs-LLM Self-Preference Bias
5.2.2 Equal Opportunity Self-Preference Bias
Similar to the LLM-vs-Human setting, we apply two complementary approaches to evaluate equal opportunity. First, we estimate conditional logistic regressions that control for content quality features. As shown in 4(b), DeepSeek-V3 exhibits the strongest and most consistent self-preference bias among the three models, showing biases of $15.5\%$ against LLaMA 3.3-70B and $12\%$ against GPT-4o. By contrast, GPT-4o does not exhibit strong self-preference relative to LLaMA 3.3-70B and, in fact, shows a negative bias when compared with DeepSeek-V3, favoring DeepSeek-V3’s outputs over its own. LLaMA 3.3-70B shows little self-preference in either comparison. Detailed conditional logistic regression results are provided in Table 4.
Second, we benchmark model decisions against human judgments of resume quality. The results from this annotation-based evaluation are more mixed. Notably, in contrast to the regression findings, LLaMA 3.3-70B exhibits a statistically significant self-preference bias of $12.92\%$ ($95\%$ CI: $12.44\%$, $13.39\%$]) when compared to GPT-4o. Meanwhile, neither DeepSeek-V3 nor GPT-4o shows significant equal opportunity self-preference, despite previously exhibiting strong statistical parity self-preference bias.
These findings highlight that LLM-vs-LLM self-preference is model-specific and generally weaker than LLM-vs-Human self-preference. The heterogeneity observed suggests that self-preference is not a uniform property of LLMs but may instead reflect model-specific factors, such as differences in stylistic alignment or the ability to recognize patterns in their own outputs.
5.3 Robustness Checks
In the following, we assess the robustness of our empirical results by considering an alternative approach to constructing counterfactual resumes in the correspondence experiment. In our baseline design, counterfactual summaries are generated from the remainder of the resume content. As a robustness check, we instead prompt each LLM to revise the existing human-written executive summary by improving its clarity and wording, while explicitly preserving the candidate’s experience and qualifications. This alternative construction mirrors a common and realistic use case of AI-assisted writing, in which job applicants employ LLMs to polish their resumes rather than generate summaries from scratch. This design allows us to examine whether AI self-preference persists under AI-assisted revision rather than reconstruction, thereby ruling out the possibility that our main results are driven by the specific method used to generate counterfactual text. Using this revised setup, we re-estimate both LLM-vs-Human and LLM-vs-LLM self-preference biases under the same evaluation framework, including conditional logistic regressions to assess equal opportunity bias.
(a) LLM-vs-Human Statistical Parity Self-preference Bias
(b) LLM-vs-Human Equal Opportunity Self-preference Bias
Figure 5: LLM-vs-Human Self-Preference Bias under Resume Revision Condition
As shown in 5(a) and 5(b), across both statistical parity and equal opportunity metrics, the qualitative patterns of LLM-vs-Human self-preference remain largely unchanged under the resume revision design. In both specifications, most evaluator LLMs continue to favor summaries revised by themselves over the original human-written summaries. Although the magnitude of self-preference is attenuated, for some models the magnitude still exceed $90\%$. Such results suggest that LLM-vs-Human self-preference is a robust phenomenon that arises even when LLM-generated text is obtained through revision rather than generation from scratch. The corresponding conditional logistic regression estimates are reported in Table 5.
Turning to LLM-vs-LLM self-preference, the revision-based design yields more heterogeneous estimates than the baseline reconstruction specification (see 6(a) and 6(b)), but several qualitative patterns persist. Consistent with the conditional logistic regression results reported in Table 6, DeepSeek-V3 continues to exhibit the strongest and most consistent equal-opportunity self-preference. Under the revision condition, DeepSeek-V3 displays equal-opportunity biases of approximately $23.2\%$ relative to GPT-4o and $8.5\%$ relative to LLaMA-3.3-70B—magnitudes that are comparable to, and in some cases larger than those observed under resume summary reconstruction. In addition, LLaMA-3.3-70B also exhibits substantial self-preference when evaluated against GPT-4o, indicating that asymmetric self-preference across model pairings persists even under the revision-based counterfactual.
(a) LLM-vs-LLM Statistical Parity Self-preference Bias
(b) LLM-vs-LLM Equal Opportunity Self-preference Bias
Figure 6: LLM-vs-LLM Self-Preference Bias under Resume Revision Condition